A 3 Story Medical building (essential facility) with ordinary masonry shear walls and ordinary reinforced concrete moment frames is to be constructed on a stiff soil base in Charleston, SC. Determine the lateral seismic forces at each floor.

Problem 11.5 A 3 story medical building (essential facility) with ordinary reinforced masonry shear walls and ordinary reinforced concrete moment frames is to be constructed on a stiff soil base in Charleston, South carolina (80° longitude, 33 N latitude). Determine the lateral seismic forces at each floor
Answer #1

Solution:-

Design seismic base shear = Ah W

Fundamental natural period = 0.075 (h)0.75

                                          Ta = 0.075 x (11.25)0.75

                                               = 0.46 second

Where h is in meter

Design horizontal seismic coefficient

Ah = Z/2 *I/R *Sa/g

    = 0.36/2 *1.5/3 *2.17 = 0.1953

Where Z = zone factor = 0.36 (for South Carolina )

For medical building( I) = 1.5

R =Response reduction factor =3

Sa/g = 1/ Ta = 1/ (0.46) = 2.17

VB =Ah* W

       = 0.1953 * 44 *105 =8.59 *105 lb

W = seismic weight of building

Design lateral seismic force at first floor

Q1 =     VB * W1 * h12/(W1h12 +W2h22 + W3h32 )

     = 8.59 *105 * 17 *105*122/(17*105* 122 + 15 * 105 * 242 + 12 * 105* 362)

     = 7.81 * 104 lb      Answer

Design lateral force at second floor

Q2 =8.59 *105 *15 *105 * 242/ ( 17 *105 *122 + 15 *105 * 242 + 12 * 105 *362 )

Q2 =2.78 * 105 lb Answer

lateral seismic force At third floor

Q3 = 8.59 * 105 * 12 * 105 * 362 / (17 *105 *122 + 15 * 105 *242 + 12 * 105 * 362)

Q3 = 5.01 * 105 lb Answer